# The Jacobi Factoring Circuit: Quantum Factoring with Near-Linear Gates and Sublinear Space and Depth

### v1

@@ 3 Factoring Squarefull Integers
$\sum_{\begin{subarray}{c}1\leq x\leq M\\ \left(\frac{x}{b}\right)=\sigma\end{subarray}}\exp\left(-\frac{2\pi ixy}{M}\right)$ $=\sum_{\begin{subarray}{c}j\in[1,B-1]\\ \left(\frac{j}{B}\right)=\sigma\end{subarray}}\sum_{\begin{subarray}{c}1\leq x\leq M\\ x\equiv j\bmod{B}\end{subarray}}\exp\left(-\frac{2\pi ixy}{M}\right)$
$=\sum_{\begin{subarray}{c}j\in[1,B-1]\\ \left(\frac{j}{B}\right)=\sigma\end{subarray}}\sum_{0\leq l\leq\left\lfloor\frac{M-j}{B}\right\rfloor}\exp\left(-\frac{2\pi i(lB+j)y}{M}\right)\text{ (writing $x=lB+j$)}$
$=\sum_{\begin{subarray}{c}j\in[1,B-1]\\ \left(\frac{j}{B}\right)=\sigma\end{subarray}}\left[\exp\left(-\frac{2\pi ijy}{M}\right)\cdot\sum_{0\leq l\leq\left\lfloor\frac{M-j}{B}\right\rfloor}\exp\left(-\frac{2\pi ilBy}{M}\right)\right]$ (3)
@@
$\sum_{\begin{subarray}{c}j\in[1,B-1]\\ \left(\frac{j}{B}\right)=\sigma\end{subarray}}\left[\exp\left(-\frac{2\pi ijy}{M}\right)\cdot\sum_{0\leq l\leq\left\lfloor\frac{M-j}{B}\right\rfloor}\exp\left(-\frac{2\pi ilBy}{M}\right)\right]$
$=$ $\sum_{\begin{subarray}{c}j\in[1,B-1]\\ \left(\frac{j}{B}\right)=\sigma\end{subarray}}\left[\exp\left(-\frac{2\pi ijy}{M}\right)\cdot\left(R+O(1)\right)\right]$
$=$ $R\cdot\left[\sum_{\begin{subarray}{c}j\in[1,B-1]\\ \left(\frac{j}{B}\right)=\sigma\end{subarray}}\exp\left(-\frac{2\pi ijy}{M}\right)\right]+O(\varphi(B))$
$=$ $R\cdot\left[\sum_{\begin{subarray}{c}j\in[1,B-1]\\ \left(\frac{j}{B}\right)=\sigma\end{subarray}}\exp\left(-2\pi ij\left(\frac{k}{B}+\delta\right)\right)\right]+O(\varphi(B))$
@@
$=$ $R\cdot\left[\left(\sum_{\begin{subarray}{c}j\in[1,B-1]\\ \left(\frac{j}{B}\right)=\sigma\end{subarray}}\exp\left(-\frac{2\pi ijk}{B}\right)\right)+O(B\varphi(B)\epsilon)\right]+O(\varphi(B))$
@@
$\left|R\cdot\left[\frac{\sqrt{B}}{2}+O(1)\right]+O(\varphi(B))\right|$
$\geq$ $\left|R\cdot\frac{\sqrt{B}}{2}\right|-O\left(\left|R\right|\right)-O(\varphi(B))$
$\geq$ $\Omega(q\sqrt{B})-O(q)-O(\varphi(B))\text{ (since $R=\Theta(q)$)}$
@@
$\Omega\left(\frac{2B}{M^{2}\varphi(B)}\cdot(q\sqrt{B})^{2}\right)=\Omega\left(\frac{q^{2}B^{2}}{M^{2}\varphi(B)}\right)\geq\Omega\left(\frac{1}{\varphi(B)}\right).$
@@ 4 Algorithm for Computing Jacobi Symbols
In this section we present one of our core technical contributions: an algorithm to compute the Jacobi symbol of $x$ mod $N$, where $N$ is classical and $x$ could be in superposition. We remark that our algorithm is also readily adaptable to computing the $\gcd$ of $x$ and $N$, much like other algorithms for computing the Jacobi symbol [Sch71, BS96, Möl08]. When $N<2^{n}$ and $x<2^{m}$, our construction requires circuit-size $\widetilde{O}(n)$ and space $\widetilde{O}(m)$, which we can exploit due to our analysis in Section 3 which allows $m$ to be smaller than $O(n)$. In contrast, the 2012 result of Li, Peng, Du and Suter [LPDS12], together with near-linear time algorithms due to [Sch71, BS96, Möl08] for computing Jacobi symbols, uses gates and space $\widetilde{O}(n)$ (see Corollary 3.2 for a formal statement of this result by [LPDS12]). Our improvements are thus along the axes of space and, as we will see in Section 4.2, depth.
We begin with an abstract algorithm (formalized in Theorem 4.1) that makes black-box use of circuits for multiplying and for computing the Jacobi symbol between equally-sized inputs. We then instantiate the circuits using explicit constructions for these subroutines [Sch71, BS96, Möl08, NZLS23] in Section 4.2.
@@ 4.1 Abstract Construction
The standard algorithms for computing the Jacobi symbol are the extended Euclidean algorithm and the binary GCD; out of the box, neither one achieves the efficiency we desire, in particular when one input is much smaller than the other. Indeed, the extended Euclidean algorithm requires space proportional to the length of the larger input, and the binary GCD’s circuit size is quadratic in the length of the larger input. Nevertheless, we can draw inspiration from an observation about the extended Euclidean algorithm: after just one iteration, the larger input is reduced to roughly the same size as the smaller one, and the entire rest of the computation has cost that scales only with the length of the smaller input. Unfortunately, it is not clear how to implement that first step—an integer division with remainder—in sublinear space. Our algorithm provides an alternative, which is both space- and depth-efficient, to this first step.
Our approach is to find a multiple $kx$ of $x$ such that both of the following are true (where we let $m\geq\lceil\log x\rceil$): (a) $N-kx$ is divisible by $2^{n-m}$; and (b) $kx<2^{n}$. Thus we can (implicitly) carry out the following chain of transformations to compute $\left(\frac{x}{N}\right)$, using the properties of the Jacobi symbol stated in Theorem 2.4:
@@
To determine such a $y=kx$, we stream through $N$ from lowest-order bit to higher-order bits, considering blocks of $m$ bits at a time. For each block, we find the appropriate multiple of $x$ to add to $y$, such that those $m$ bits of $y$ match with the corresponding bits of $N$ (and thus the corresponding bits of $N-y$ are zero). At each step, we compute the needed multiplicative factor in an $m$-bit control register $\mathsf{ctrl}$; crucially, we show that $\mathsf{ctrl}$ is efficient to both compute and uncompute (see our analyis of Algorithm 4.2).
The fact that $\mathsf{ctrl}$ can be uncomputed efficiently is key to our space savings, as it means each individual iteration of our algorithm is reversible and no garbage bits accumulate throughout the iterations. Furthermore, once the current block of $m$ bits matches those of $N$, these bits will not depend on anything quantum (since $N$ is classical). The qubits representing them are thus in a classically-known state, and unentangled from all other qubits in the computation. We use this fact to recycle these qubits. Ultimately, we only ever store the leading $O(m)$ bits of the partial value $y$ that is being computed. In Algorithm 4.2, we denote the register holding this sliding window of $O(m)$ bits as $z$.
In this section, we first describe how to utilize the value $kx$ to compute the Jacobi symbol (Algorithm 4.1), then present our procedure for obtaining the value $kx$ (Algorithm 4.2), and then finally prove our claimed performance guarantees.
@@
    Compute $\mathsf{out}\leftarrow\mathsf{out}\cdot\left(\frac{s}{x^{\prime}}\right)$, where the Jacobi symbol $\left(\frac{s}{x^{\prime}}\right)$ is computed via the algorithms of [Sch71, BS96, Möl08], made reversible via standard techniques [Ben73, Ben89, LS90].
@@
We proceed by induction on $j$. For the base case where $j=0$, we have $z_{0}=0\Rightarrow y_{0}=0$. All three conditions are now evident. Now, for the inductive step, assume that we have shown the result for some $j\geq 0$ and wish to show the result for $j+1$. We will examine the execution of iteration $j$ of step 3 of Algorithm 4.2 to complete the induction. Firstly, by definition we have the following:
@@
and Equation (6) can also be written as:
$2^{(j+1)m}z_{j+1}+\left(N\bmod 2^{(j+1)m}\right)$ $\equiv N\pmod{2^{(j+1)m}}$
$\Leftrightarrow y_{j+1}$ $\equiv N\pmod{2^{(j+1)m}},$
thus establishing condition 1. Now, we also have that:
@@
We proceed by showing that each step of Algorithm 4.1 can be implemented reversibly with the stated complexities. We note that $S_{\mathrm{Jac}}(m)$ must be at least linear in $m$ because the Jacobi symbol depends on the entire input; therefore our space complexity is lower bounded by $O(m)$.
@@ 4.2 Implications: Factoring Certain Integers in Sublinear Space and Depth
In this section, we instantiate Algorithms 4.1 and 4.2. Here we focus on asymptotic costs, leaving the optimization of circuits for practical problem sizes to future work. For multiplication, we use a parallelized quantum circuit for Schönhage-Strassen multiplication [SS71], by which the product of two $t$-bit quantum integers can be computed in gate count $\widetilde{O}(t)$ and depth $\mathsf{polylog}(t)$, using $\widetilde{O}(t)$ total qubits [NZLS23]. For computing the Jacobi symbol of two inputs of size $t$, there exist classical algorithms with complexity $\widetilde{O}(t)$ [Sch71, BS96, Möl08]; by standard reversible circuit techniques these algorithms can be made into quantum circuits with gate count, depth, and qubit count all at most $\widetilde{O}(t)$ [Ben73, Ben89, LS90]. The following corollary results directly from instantiating Algorithms 4.1 and 4.2 with these constructions.
@@
Remark (Near-optimal parallelism for small $Q$).
@@
We now state the implications of Corollary 4.7 when factoring numbers of the form $N=P^{2}Q$ with $\log Q=\widetilde{\Theta}((\log N)^{2/3})$ and $P,Q$ are prime. As summarized in Section 2.3, there are no known classical special-purpose factoring algorithms that perform better than the general number field sieve [Pol93, LLMP90, BLP93] on integers of this form. Yet by Corollary 4.7, it is possible to quantumly factor these integers in much less space and depth than would be required for generic integers [Sho97, Reg23] or even generic squarefull integers [LPDS12]. Indeed, applying Corollary 4.7 to numbers of that form yields the following result:
@@
5 Fully Factoring Integers with Distinct Exponents in their Prime Factorization
Here, we provide a black-box reduction that shows that any algorithm achieving the guarantees of Theorem 3.1 can in fact be used to fully factor integers of the form $N=p_{1}^{\alpha_{1}}\ldots p_{r}^{\alpha_{r}}$ with $\alpha_{1},\ldots,\alpha_{r}$ positive and distinct.

### v3

@@ 3 Factoring Squarefull Integers
$\sum_{1\leq x\leq M}\left(\frac{x}{B}\right)\exp\left(-\frac{2\pi ixy}{M}\right)$ $=\sum_{1\leq j\leq B-1}\sum_{\begin{subarray}{c}1\leq x\leq M\\ x\equiv j\bmod{B}\end{subarray}}\left(\frac{j}{B}\right)\exp\left(-\frac{2\pi ixy}{M}\right)$
$=\sum_{1\leq j\leq B-1}\sum_{0\leq l\leq\left\lfloor\frac{M-j}{B}\right\rfloor}\left(\frac{j}{B}\right)\exp\left(-\frac{2\pi i(lB+j)y}{M}\right)\text{ (writing $x=lB+j$)}$
$=\sum_{j\in[1,B-1]}\left[\left(\frac{j}{B}\right)\exp\left(-\frac{2\pi ijy}{M}\right)\cdot\sum_{0\leq l\leq\left\lfloor\frac{M-j}{B}\right\rfloor}\exp\left(-\frac{2\pi ilBy}{M}\right)\right].$ (3)
@@
$\sum_{j\in[1,B-1]}\left[\left(\frac{j}{B}\right)\exp\left(-\frac{2\pi ijy}{M}\right)\cdot\sum_{0\leq l\leq\left\lfloor\frac{M-j}{B}\right\rfloor}\exp\left(-\frac{2\pi ilBy}{M}\right)\right]$
$=$ $\sum_{j\in[1,B-1]}\left[\left(\frac{j}{B}\right)\exp\left(-\frac{2\pi ijy}{M}\right)\cdot\left(R+O(1)\right)\right]$
$=$ $R\cdot\left[\sum_{j\in[1,B-1]}\left(\frac{j}{B}\right)\exp\left(-\frac{2\pi ijy}{M}\right)\right]+O(\varphi(B))$
$=$ $R\cdot\left[\sum_{j\in[1,B-1]}\left(\frac{j}{B}\right)\exp\left(-2\pi ij\left(\frac{k}{B}+\delta\right)\right)\right]+O(\varphi(B))$
@@
$=$ $R\cdot\left[\left(\sum_{j\in[1,B-1]}\left(\frac{j}{B}\right)\exp\left(-\frac{2\pi ijk}{B}\right)\right)+O(B\varphi(B)\epsilon)\right]+O(\varphi(B))$
@@
$\left|R\cdot\left[\sqrt{B}+O(1)\right]+O(\varphi(B))\right|$
$\geq$ $\left|R\cdot\sqrt{B}\right|-O\left(\left|R\right|\right)-O(\varphi(B))$
$\geq$ $\Omega(q\sqrt{B})-O(q)-O(\varphi(B))\text{ (since $|R|=\Theta(q)$)}$
@@
$\Omega\left(\frac{B}{M^{2}\varphi(B)}\cdot(q\sqrt{B})^{2}\right)=\Omega\left(\frac{q^{2}B^{2}}{M^{2}\varphi(B)}\right)\geq\Omega\left(\frac{1}{\varphi(B)}\right).$
@@
Remark 3 (On the use of Gauss sums).
In this analysis, we made use of the “strong half” of Lemma 2.16, namely its conclusion in the case where $\gcd(k,m)=1$. On the other hand, the original analysis by [LPDS12] only makes use of the “weak half” i.e. the case where $\gcd(k,m)>1$. The reason we use Gauss sums in all their power is because of our adaptation of Shor’s analysis [Sho97] to the setting where the initial superposition ranges only up to $\mathsf{poly}(Q)$. As in Shor’s analysis, we lower bounded the amplitude of the final (post-QFT) state on each “useful” value and this requires a strong Gauss sum bound.
If we wanted to get around the need for the strong Gauss sum bound, we could instead adapt the tighter analysis of Regev’s factoring algorithm [Reg25] based on the Poisson summation formula. This would only require the “weak” Gauss sum result (as in [LPDS12]), but would likely require us to start with a discrete Gaussian superposition instead of a uniform superposition. Given the concrete overheads involved in preparing discrete Gaussian superpositions [GR02, Reg09] and our interest in the potential practicality of our algorithms, we chose to present the simpler algorithm with just a uniform superposition.
We thank Oded Regev for pointing out to us that the use of “strong” Gauss sum bounds in our analysis could be circumvented.
@@ 4 Algorithm for Computing Jacobi Symbols
In this section we present one of our core technical contributions: an algorithm to compute the Jacobi symbol of $x$ mod $N$, where $N$ is classical and $x$ could be in superposition. We remark that our algorithm is also readily adaptable to computing the $\gcd$ of $x$ and $N$, much like other algorithms for computing the Jacobi symbol [Sch71, TY90, BS96, Möl08]. When $N<2^{n}$ and $x<2^{m}$, our construction requires circuit-size $\widetilde{O}(n)$ and space $\widetilde{O}(m)$, which we can exploit due to our analysis in Section 3 which allows us to restrict $m\ll n$. In contrast, the 2012 result of Li, Peng, Du and Suter [LPDS12], together with near-linear time algorithms due to [Sch71, TY90, BS96, Möl08] for computing Jacobi symbols, uses gates and space $\widetilde{O}(n)$ (see Corollary 3.2 for a formal statement of this result by [LPDS12]). Our improvements are thus along the axes of space and, as we will see in Section 4.2, depth.
We begin with an abstract algorithm (formalized in Theorem 4.1) that makes black-box use of circuits for multiplying and for computing the Jacobi symbol between equally-sized inputs. We then instantiate the circuits using explicit constructions for these subroutines [Sch71, TY90, BS96, Möl08, NZLS23] in Section 4.2.
@@ 4.1 Abstract Construction
The standard algorithms for computing the Jacobi symbol are the extended Euclidean algorithm and the binary GCD; out of the box, neither one achieves the efficiency we desire, in particular when one input is much smaller than the other. Indeed, existing circuits for both algorithms require space proportional to the length of the larger input. Nevertheless, we can draw inspiration from an observation about the extended Euclidean algorithm: after just one iteration, the larger input is reduced to roughly the same size as the smaller one, and the entire rest of the computation has cost that scales only with the length of the smaller input. Our task thus boils down to “just” computing $N\bmod{x}$ in sublinear space. The naive idea would be to use standard long division or a variant thereof. However, while long division only needs to look at $O(\log(x))$ bits of $N$ at a time, we have to keep track of the intermediate values to make the computation reversible. Sadly, keeping track of these excess values appears to require linear space.
Our solution, which is inspired by the binary GCD algorithm [BS96] and strongly resembles Montgomery reduction [Mon85], is to perform a “flipped” variant of long division. In the Euclidean algorithm, the computation of $N\bmod x$ can be thought of as a way to find a multiple $\kappa x$ such that $N-\kappa x<x$ — in particular, all but the lowest $m$ bits of $N-\kappa x$ are zero. Long division achieves this by starting from the most significant bit (MSB) and iterating towards the least significant bit (LSB), zeroing out bits along the way. We flip this idea on its head: we compute a multiple $kx$ such that all but the highest-order $m$ bits of $N-kx$ are zero—by starting from the LSB and iterating towards the MSB. From here, there are three key ideas:
  1. 1.
    The value $kx$ can be built up bit by bit via a loop in which each iteration is entirely reversible. We can easily uncompute intermediate states by performing a simple comparison that depends only on the MSBs of our current state.
  2. 2.
    Furthermore, once a lower-order bit of $kx$ matches that of $N$, this bit will not depend on anything quantum (since $N$ is classical). This qubit is thus in a classically-known state, and unentangled from all other qubits in the computation. We can use this fact to recycle lower-order qubits as we set them. Ultimately, we only ever store the leading $O(m)$ bits of the partial value $y=kx$ that is being computed. In Algorithm 4.2, we will denote the register holding this sliding window of $O(m)$ bits as $z$.
  3. 3.
    The above two ideas on their own already suffice to obtain a circuit that computes the Jacobi symbol in sublinear space $O(m)$. However, its gate count will be $O(nm)$. This is because, when setting one bit of $k$ at a time, we will ultimately have to carry out “schoolbook” arithmetic comprising additions.
    Instead, we can set $k$ in batches of $m$ bits each. This allows us to benefit from fast integer multiplication algorithms [SS71, HvdH21, KMY24], and will ultimately drop the gate count to $\widetilde{O}(n)$, which is near-linear.
At the end, our quantum computer will hold the value $(N-kx)/2^{n-m}$, where $kx$ is a multiple such that both of the following are true (where we let $m\geq\lceil\log x\rceil$): (a) $N-kx$ is divisible by $2^{n-m}$; and (b) $kx<2^{n}$. We can now proceed from here in one of two ways:
  • This actually suffices to complete the first step of extended Euclidean and compute $N\bmod{x}$, since we have:
    $N$ $\equiv\frac{N-kx}{2^{n-m}}\cdot 2^{n-m}\pmod{x},$
    and $2^{n-m}\bmod{x}$ is very efficiently computable. This is asymptotically a satisfactory solution, but adds unnecessary indirection and concrete efficiency to our overall algorithm, as we will see next.
    We thank Daniel J. Bernstein for pointing out this variant of our algorithm to us.
  • Rather than taking the extra step as above to compute $N\bmod{x}$, we could just start from $(N-kx)/2^{n-m}$ and directly carry out the following chain of transformations to compute $\left(\frac{x}{N}\right)$, using the properties of the Jacobi symbol stated in Theorem 2.4:
@@
    This is the approach that we will take, as detailed in Algorithm 4.1.
We first describe how to utilize the value $kx$ to compute the Jacobi symbol (Algorithm 4.1), then present our procedure for obtaining the value $kx$ (Algorithm 4.2), and then finally prove our claimed performance guarantees.
@@
    Compute $\mathsf{out}\leftarrow\mathsf{out}\cdot\left(\frac{s}{x^{\prime}}\right)$, where the Jacobi symbol $\left(\frac{s}{x^{\prime}}\right)$ is computed via the algorithms of [Sch71, TY90, BS96, Möl08], made reversible via standard techniques [Ben73, Ben89, LS90].
@@
Data: Efficiency parameters $m,n$ such that $m|n$, and positive integers $N<2^{n}$ and $x<2^{m}$ with $x$ odd
@@
Remark 4.
We state our result in terms of explicitly writing down the Jacobi symbol for the purposes of clarity and to emphasize that this also works equally well in the classical reversible setting. However, as detailed in Section 3, our algorithms will ultimately apply the Jacobi symbol in the phase, thus computing the unitary |$\ket{x}\mapsto j_{N}(x)\ket{x}$ for some function $j_{N}(x)$ that is equal to the Jacobi symbol whenever $\gcd(x,N)=1$. This can easily be achieved either by using the above algorithm out of the box and applying a phase kickback [CEMM98], or by going through Algorithms 4.1 and 4.2 and modifying them to write the output in the phase rather than in a single-qubit register.
@@
Item 1 is straightforward. To establish the other two items, we proceed by induction on $j$. For the base case where $j=0$, we have $z_{0}=0\Rightarrow y_{0}=0$. All three conditions are now evident. Now, for the inductive step, assume that we have shown the result for some $j\geq 0$ and wish to show the result for $j+1$. We will examine the execution of iteration $j$ of step 3 of Algorithm 4.2 to complete the induction. Firstly, by definition we have the following:
@@
Now, we also have that:
@@
We proceed by showing that each step of Algorithm 4.1 can be implemented reversibly with the stated complexities. We note that $S_{\mathrm{Jac}}(m)$ must be at least linear in $m$ because the Jacobi symbol depends on the entire input; therefore our space complexity is lower bounded by $\Omega(m)$.
@@ 4.2 Implications: Factoring Certain Integers in Sublinear Space and Depth
In this section, we instantiate Algorithms 4.1 and 4.2. Here we focus on asymptotic costs, leaving the optimization of circuits for practical problem sizes to future work. For multiplication, we use a parallelized quantum circuit for Schönhage-Strassen multiplication [SS71], by which the product of two $t$-bit quantum integers can be computed in gate count $\widetilde{O}(t)$ and depth $\mathsf{polylog}(t)$, using $\widetilde{O}(t)$ total qubits [NZLS23]. For computing the Jacobi symbol of two inputs of size $t$, there exist classical algorithms with complexity $\widetilde{O}(t)$ [Sch71, TY90, BS96, Möl08]; by standard reversible circuit techniques these algorithms can be made into quantum circuits with gate count, depth, and qubit count all at most $\widetilde{O}(t)$ [Ben73, Ben89, LS90]. The following corollary results directly from instantiating Algorithms 4.1 and 4.2 with these constructions.
@@
Remark 5 (Near-optimal parallelism for small $Q$).
@@
We now state the implications of Corollary 4.7 when factoring numbers of the form $N=P^{2}Q$ with $\log Q=\widetilde{\Theta}((\log N)^{2/3})$ and $P,Q$ are prime. As summarized in Section 2.3, there are no known classical special-purpose factoring algorithms that perform better than the general number field sieve [Pol93, LLMP90, BLP93] on integers of this form. Yet by Corollary 4.7, it is possible to quantumly factor these integers in much less space and depth than would be required for generic integers [Sho97, Reg25] or even generic squarefull integers [LPDS12]. Indeed, applying Corollary 4.7 to numbers of that form yields the following result:
@@
5 Completely Factoring Integers with Distinct Exponents in their Prime Factorization
Here, we provide a black-box reduction that shows that any algorithm achieving the guarantees of Theorem 3.1 can in fact be used to completely factor integers of the form $N=p_{1}^{\alpha_{1}}\ldots p_{r}^{\alpha_{r}}$ with $\alpha_{1},\ldots,\alpha_{r}$ positive and distinct.